A silica optical fibre has a core refractive index of 1.50 and a cladding refractive index of 1.45. Determine the critical angle at the core cladding interface. *
1
Expert's answer
2020-12-08T10:51:49-0500

Core refractive index "n_1 = 1.50"

Cladding refractive index "n_2 = 1.45"

Snell’s Law:

"sinθ_c = \frac{n_2}{n_1}"

"\u03b8_c" is the critical angle

Critical angle at the core cladding interface:

"θ_c = sin^{-1}(\frac{n_2}{n_1}) \\

= sin^{-1}(\frac{1.45}{1.50}) \\

= 75.16 º"

Answer: 75.16 º